1/7 - 10/7e + 3/e^2 divided by 1/7 - 1/e
Answer by stanbon(75887)

1/7 - 10/7e + 3/e^2 divided by 1/7 - 1/e
-----------------------
(1/e) = w
-------------
3w^2 -(10/7)w + (1/7) = 0
21w^2 - 10w + 1 = 0
(3w-1)(7w-1) = 0
w = 1/3 or w = 1/7
-------------------------
substitute back to "e"
1/e = 1/3 or 1/e = 1/7
e = 3 or e = 7
====================
Cheers,
Stan H.